OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). 3x + 2y = 2 -2x + y = 8

OpenStudy (anonymous):

frm second eqn y = 8 + 2x substitute in first eqn 3x + 2(8 + 2x) = 2 3x + 16 + 4x = 2 7x = 2 - 16 x = -14/7 x = -2

OpenStudy (anonymous):

now substitute x = -2 in second eqn -2(-2) + y = 8 4 + y = 8 y = 8 - 4 y = 4 so solution is (-2,4)